OpenStudy (anonymous):

What is the range of the data? Chart: temperature 75 80 95 65 70 95 85 80 65 70 85 95 choices: 40 30 20 65

OpenStudy (anonymous):

range=subtract highest value with lowest value.

OpenStudy (anonymous):

30. Range = biggest value - smallest value

OpenStudy (callisto):

range = largest - smallest = 95 - 65 = ?

OpenStudy (anonymous):

30

OpenStudy (callisto):

Yes

OpenStudy (anonymous):

95-65=30.

OpenStudy (anonymous):

thanks :) <3

OpenStudy (anonymous):

<3

OpenStudy (anonymous):

np.

OpenStudy (callisto):

welcome